What is the thermal diffusivity (in m³/s) of the material having properties, p-7500 kg/m³, Cp 400 J/(kg-K), k-15 W/(m-K).?


1
Expert's answer
2021-08-20T01:34:14-0400

Thermal diffusivity=(Thermal Conductivity)/(density)(spe cific heat at constant pressures).

Thermal diffusivity = 15/7500*400

=0.5×10-5W.m2/J

Watt is J/s

Thermal diffusivity=0.5×10-5m2/s